From: NASA HQ/SpaceRef/NASA Watch
Posted: Thursday, February 10, 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
Early before breakfast and first exercise, FE Salizhan Sharipov and CDR/SO Leroy Chiao completed their 9th session with the Russian crew health-monitoring program's medical assessment MO-9/Biochemical Urinalysis. Afterwards, the FE stowed the hardware. [MO-9 is conducted regularly every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for US crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the "PHS/Without Blood Labs" exam. The analysis uses the sophisticated in-vitro diagnostic apparatus Urolux developed originally for the Mir program. Afterwards, the data are entered in the medical equipment computer (MEC)'s special IFEP software (In-Flight Examination Program).]
Later, Sharipov took his 4th session with the cardiological experiment PZEh MO-1 (Study of the Bioelectric Activity of the Heart at Rest), with Chiao assisting as CMO. [During the 30-min. test, the crew tagged up with ground specialists on a Russian ground site (RGS) pass on Daily Orbit 5 (3:40am EST) via VHF and downlinked data from the Gamma-1M ECG (electrocardiograph) for about 5-6 minutes.]
With the Elektron O2 generator remaining off, the FE worked on the Russian harmful impurities removal system (BMP), continuing the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. [Regeneration of each of the two cartridges takes about 24 hours. Elektron was deactivated on 2/7 after ppO2 (oxygen partial pressure) approached the upper concentration limit of 24.1% set by Flight Rule as a flammability safety margin.]
Ending the decontamination period for the two new CSA-CPs (Compound Specific Analyzer-Combustion Products) #1018 ϻ that arrived with Progress 16, Chiao changed the battery in #1018 and zero-calibrated both units. [The two old instruments, #1015 & #1016, were stowed, and #1018 & #1019 are now prime and backup, respectively.]
Afterwards, Leroy used the new CSA-CPs for the monthly cabin air spot check, taking readings for O2 (oxygen), CO (carbon monoxide), HCN (hydrogen cyanide), HCl (hydrogen chloride) in SM and Lab, as well as battery ticks for calldown.
The crew performed the periodic comprehensive cabin air and surface sample collections for bacterial and fungal analysis, starting off with Chiao employing the MAS (Microbial Air Sampler) kit to collect air samples in Lab, Node and Service Module (SM), and the SSK (Surface Sample Kit) in Lab, Node and FGB. (Last time done: 11/5/04) [In the FGB, three samples were collected at specific panels showing stained areas. Analysis of previous samples (11/5/04) from two FGB panels has confirmed fungal contamination. Today's sampling of these panels will evaluate the effectiveness of the disinfection activities that have been performed since. MAS and SSK samplings are done once per month for the first three months, and once every three months thereafter. Bacterial and fungal air samples are taken at two locations in each module. The colony growth on the SSK sampling slides and MAS Petri dishes will be analyzed after five days of incubation. For the onboard visual analysis of SSK media slides and MAS Petri dishes, the crew has a procedure for visual inspection of samples for bacterial and fungal colony growths after appropriate incubation periods.]
Air samples were also collected by Chiao with the new Dual Sorbent Tube (DST), instead of the old SSAS (Solid Sorbent Air Sampler), in the center of the Lab and SM. (Last time done: 1/7).
FE Sharipov used the Russian AK-1M sampler in the SM and FGB for air sampling and also checked the SM for Freon with the AK-1M-F. Then, looking for CO (carbon monoxide), he took air samples in the SM with the IPD-CO Draeger tubes sampler. (Last time done: 1/7)
Continuing the current outfitting of the Russian ASN-M satellite navigation system in the SM for the European ATV (Automated Transfer Vehicle), extending through 2/11, the FE today connected the ASN-M to the SUBA onboard equipment control system. The work was again supported by S-band tagup with ground specialists. The goal is to complete all of the internal ATV GLONASS (the Russian GPS equivalent) hardware and cable installation activities by 2/16 to support Russian segment (RS) GPS testing. [In tomorrow's ASN activity, Salizhan will connect the ASN-M to the SBI onboard measuring system, completing the installation. Although not yet operational, ASN-M uses GLONASS satellites to update the ISS state vector (SV, position & velocity plus time) without using the ground (which currently has to uplink daily SV updates) or requiring SV transfers from the U.S. segment (USOS) from time to time. The ASN equipment was originally factory-installed in the SM but was found faulty and had to be returned to the ground. After repair it was shipped again to the station on Progress 11 and re-installed by Yuri Malenchenko on 7/8/03, followed by various troubleshooting attempts and the current ATV mods.]
CDR Chiao performed the periodic filter cleaning on the autonomous PCG-STES010 (Protein Crystal Growth-Single Locker Thermal Enclosure System) payload in the Lab. [For the PCG-STES, today is the 806th day of continuous powered operation onboard the ISS, which is unprecedented for this experiment hardware. Its previous run time onboard Mir did not exceed 200 days. Nominally, PCG-STES010 powered operations are expected to continue while plans are finalized for its return on LF-1 (STS-114) later this year.]
Sharipov completed the regular periodic download of data & imagery collected of the BIO-5 Rasteniya-2 ("Plants-2") experiment to the computer for subsequent downlink to the ground. Working off the Russian discretionary job jar, he then completed the regular daily inspection of the Lada-5 greenhouse equipment, including filling its water canister as required.
Also from the "job jar" list, the FE was to downlink a protocol for assessing issues with the Russian "Sigma" application, a ballistic navigation program to compute the station's ground track on the Earth.
Chiao printed out new pages with SODF (Station Operations Data File) procedures, such as for upcoming GPS activation ops, i.e., authorization & prioritizing of GPS in the G&NC (guidance, navigation & control) software, powering on and initialization.
The CDR also downlinked the ADUM OPE (Advanced Diagnostic Ultrasound in Microgravity/On-board Proficiency Enhancer) file via OCA comm, which could not be transmitted after yesterday's OPE session.
In preparation for tomorrow's planned SSRMS (Space Station Remote Manipulator System) proficiency exercise, CDR Chiao connected the UOP DCP (utility outlet panel/display & control panel) bypass power cable at the Lab RWS (robotics work station), then reviewed the current version of the DOUG (Dynamic Operational Ubiquitous Graphics) software. [Used during Robotics/SSRMS operations, DOUG is a software program on the MSS (mobile service system) RWS laptops that provides a birdseye-view graphical image of the external station configuration and the SSRMS arm, showing its real-time location and configuration on the display during its operation. The objectives of tomorrow's Robotics ops are for the crew to maintain operator proficiency and perform engineering tests to troubleshoot the "sticky grapple fixture" problem, then stow the SSRMS in its EVA-12 viewing position. The testing is a repeat of the testing performed by the crew on 12/14/04. The task involves maneuvering the arm to the Lab FRGF (flight releasable grapple fixture), followed by pre-grapple and manual grappling of the fixture. The FRGF will then be released and the end effector backed out using a special push-off and dither technique with a limped wrist joint to counteract the loads-and-friction-induced sticking phenomenon encountered earlier. Once the grapple pin is cleared, MCC-Houston will evaluate the results of the release test and the allocated time remaining to determine if there was sufficient time for an additional capture/release cycle.]
The CDR conducted the routine daily maintenance of the SOZh environment control & life support system in the SM, including the weekly inspection of the air/liquid condensate separator apparatus (BRPK). The FE meanwhile prepared the regular IMS (Inventory Management System) "delta" file for the daily automated export/import to the three IMS databases on the ground. [The BRPK inspection is mostly concerned with the condensate in a CWC (contingency water container) and the EDV with water transferred from the KTV nonpotable water, to check for presence of air bubbles by removing the outer container cover and looking through the backlit transparent shell.]
Leroy did his daily checkup of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPC's temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]
The crew worked out in accordance with their daily 2.5-hr. physical exercise protocol on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic). Salizhan's daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 3 of a new set) and one hour on VELO.
Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
At ~12:54pm EST, Leroy set up the Sputnik-SM Kenwood D700 amateur radio station in the SM and conducted a 10-min. ham radio session with students at Fairview Elementary School in Mount Prospect, IL (a "Leroy Chiao crew pick"). [Mount Prospect, IL, is located 22 mi. NW of downtown Chicago.]
A joint US/Russia investigation of the recent EVA-12 has determined that at least one crewmember inadvertently entered a designated keep-out zone (KOZ) during SM roll thruster activity (for CMG desaturation). Safety procedures are being modified on both sides to ensure future prevention of KOZ violation during spacewalks. No traces of contamination on the Orlan-M suits were found after the EVA-12 incident. [SM thruster keep-out zones are imaginary 60-deg. cones (i.e., with 120 deg. central angle) with apex at the roll thruster nozzles. Pitch and yaw are controlled by Progress thrusters.]
Today's CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Mekong River Delta, Vietnam (weather was predicted to be clear for general mapping. Overlapping image series of the delta are useful for tracking channel and vegetation changes. Changes in land use both upstream and within the delta itself can significantly [and rapidly] alter the hydrology, geomorphology, and ecology of the delta), Internal Waves, North Atlantic (the Iberian Peninsula coastline remains mostly cloud-free, providing an opportunity for internal wave photography. Looking for the sunglint point to the right of track off the northwestern tip of the Peninsula [westernmost Spain]), and Lower Amazon River Basin, S. America (this overpass took the station across the full width of the delta. Overlapping image swaths of the feeder channels and islands are useful for monitoring hydrologic and geomorphic changes to the delta. These changes are primarily due to land use change [forest clearing, soil loss, etc.] within the Amazon River watershed).
CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 10 crew visit:
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
ISS Orbit (as of last night, 3:50pm EST [= epoch]):
// end //